OpenStudy (anonymous):

Find the derivative of each function defined as follows: y=5x^-5-6x^-2+13x^-1

OpenStudy (imstuck):

You are just differentiating x with respect to y?

OpenStudy (imstuck):

Do you know how the power rule works in derivatives?

OpenStudy (anonymous):

derivative of x^(-n) is -n*x^(-n-1)

OpenStudy (anonymous):

Hi! yes, I know how the power rule works

OpenStudy (imstuck):

Look at what matricked did.

OpenStudy (anonymous):

so derivative of 5x^-5 = -5*5x^-6 = -25 x^-6 and similar for the rest..

OpenStudy (imstuck):

\[5x ^{-5}=-5x ^{-5-1}=-5x ^{-6}\]

OpenStudy (imstuck):

That's just for the first term, but the others are done the same way.

OpenStudy (anonymous):

Ok, let me try it out. Thanks!

OpenStudy (anonymous):

My answer is: \[25x ^{-6}+12x^{-3}+13x ^{-2}\]

OpenStudy (anonymous):

Use the power rule and the signal rule:\[y=x^k\\y'=kx^{k-1}\]_______________________________________________________________ \[y=5x^{-5}-6x^{-2}+13x^{-1}\]\[y'=5*(-5)*x^{-5-1}-6*(-2)*x^{-2-1}+13*(-1)*x^{-1-1}\]\[y'=-25x^{-6}+12x^{-3}-13x^{-2}\]

OpenStudy (anonymous):

Do I have to do anything about the negative exponents? or can this be my final answer?

OpenStudy (anonymous):

You forgot the signal rule, if the exponents is negative, you should multiple the rest by the negative exponent, for example:\[y=4x^{-3}\]The derivative will be (pay attention in the exponent\[y'=4*(-3)*x^{-3-1}\]the answer will be\[y'=-12*x^{-4}\]The second derivative will be\[y''=-12*(-4)*x^{-4-1}\]\[y''=48*x^{-5}\]

OpenStudy (anonymous):

Ok, thank you!

OpenStudy (anonymous):

You're welcome ;)